OpenStudy (pphalke):

find the interest and the total amount for $825 invested for 10 years at 4% APR compounded annually

OpenStudy (photon336):

Start off with the compound interest formula \[A_{0}(1+\frac{ r }{ n })^{nt} = A \]

OpenStudy (photon336):

Where A0 is your initial investment

OpenStudy (pphalke):

is there a compound interest calculator

OpenStudy (photon336):

\[825*(1+\frac{ .04 }{ 1 })^{10} = ?\]

OpenStudy (photon336):

@pphalke what did you get as an answer?